# Explanation Evaluation

### v1

@@ Abstract
Evaluating free-text explanations is a multifaceted, subjective, and labor-intensive task. Large language models (LLMs) present an appealing alternative due to their potential for consistency, scalability, and cost-efficiency. In this work, we present [Uncaptioned image]Acorn, a new dataset of 3,500 free-text explanations and aspect-wise quality ratings, and use it to gain insights into how LLMs evaluate explanations. We observed that replacing one of the human ratings sometimes maintained, but more often lowered the inter-annotator agreement across different settings and quality aspects, suggesting that their judgments are not always consistent with human raters. We further quantified this difference by comparing the correlation between LLM-generated ratings with majority-voted human ratings across different quality aspects. With the best system, Spearman’s rank correlation ranged between 0.53 to 0.95, averaging 0.72 across aspects, indicating moderately high but imperfect alignment. Finally, we considered the alternative of using an LLM as an additional rater when human raters are scarce, and measured the correlation between majority-voted labels with a limited human pool and LLMs as an additional rater, compared to the original gold labels. While GPT-4 improved the outcome when there were only two human raters, in all other observed cases, LLMs were neutral to detrimental when there were three or more human raters. We publicly release the dataset to support future improvements in LLM-in-the-loop evaluation. [^1] https://github.com/a-brassard/ACORN
@@ 1 Introduction
Natural language processing systems that not only generate correct output, but also provide an explanation (Miller, 2017) of why that particular output is correct, are desirable for several reasons, such as increasing trustworthiness (Floridi, 2019), compliance with “right to explanation” laws (e.g., European Parliament and Council of the European Union, ), increasing interpretability (Jacovi and Goldberg, 2020, but cf. Lipton, 2018), as well as system improvement and knowledge discovery (Adadi and Berrada, 2018). However, this immediately raises the question of how to evaluate system-generated explanations in an efficient and effective manner.
Since explanations are typically free-form text, automatic evaluation of explanations suffers the well-known, but as of yet unresolved, weaknesses of automatic evaluation measures (Celikyilmaz et al., 2021), while human evaluation is characterized by low scalability, high costs, subjectivity, and inconsistency (Hartmann and Sonntag, 2022). LLM-based evaluation presents an appealing alternative due to its potential high scalability, low cost, and consistency.
@@
Specifically, we first considered whether LLMs’ labels deviate from what would be expected from human raters. We compared inter-annotator agreement between all-human annotation (five raters) and when one of the raters is replaced by an LLM. We found that most LLMs reduced the agreement, indicating that their judgments skew outside of expected human rating.
GPT-4, however, had consistently less impact on the agreement, suggesting that it may still have potential as a replacement for human judgments. To verify this, we considered two scenarios: one where the LLM replaces the full human evaluation, and one where it is used only as an additional rater. In other words, we compared LLMs to humans collectively and individually.
With the best-performing model, Spearman’s rank correlation between majority-voted gold labels and LLM-generated ones ranged between 0.52 and 0.95, depending on the aspect, averaging 0.72. This indicates that the LLM’s labels are not entirely in line with human judgments, but they are not entirely random either, and their usefulness may depend on the specific evaluation criteria and usage case.
As for using the LLM as a single rater, we considered whether, in a limited-budget scenario, it would be beneficial to use an LLM as an additional rater. Specifically, we compared whether the majority-voted labels of a reduced set of raters and with an added LLM had a higher correlation with the original majority-voted labels. In most cases, the addition of LLMs either had no effect or was detrimental, suggesting that they are not a reliable replacement for human raters. One exception was GPT-4, which improved correlation in a scenario with only two human raters. However, the resulting labels are still less reliable than those from three or more human raters.
In summary, we quantified the consequences of using LLMs as a replacement or addition to human evaluation, concluding that most have rating biases that diverge from human consensus, leading to different outcomes. GPT-4 had a beneficial effect when there were only two other human raters, but did not compensate as much as an additional human rater did. We release [Uncaptioned image]Acorn as a benchmark for future research on automatic evaluation of explanations.
@@ 2 Building [Uncaptioned image]Acorn: Evaluation Criteria and Data Sources
In a typical commonsense reasoning test, a model selects the most plausible answer for a multiple-choice question. In a predict-and-explain setting, the model additionally generates an explanation to justify the selected answer, where we encounter the challenge of evaluating these explanations. Thus, we first define a set of rating criteria (§2.1) and collect human ratings for a selection of existing and newly collected textual explanations (§2.2, §2.3).
@@ 2.1 Rating Criteria
Table 1: Explanation rating criteria used in this study. The first two criteria target faithfulness and general explanation quality, while the rest covers specific quality aspects.
As of now, there is no consensus on which criteria to use when evaluating explanations. Our selection was informed by the following. Explanation evaluation falls under the umbrella of natural language generation (NLG) evaluation. We boiled down commonly used criteria (Howcroft et al., 2020) to two overarching aspects: surface-level criteria and information/content-level criteria, namely fluency and information sufficiency or minimality. Additionally, we introduced criteria informed by known challenges of free-text explanations: an LLM-generated explanation is not necessarily faithful to a model’s inner workings (faithfulness; Lipton, 2018), and can include hallucinated untrue facts (factuality; Rawte et al., 2023). Finally, inspired by Miller (2017), we add a criterion aligned with an ideal explanation’s structure (contrastiveness) and a retain a non-specific overall rating intended to capture potentially informative or useful aspects that we have not explicitly covered. Our resulting criteria are summarized in Table 1. During crowdsourcing, workers were given detailed guidelines and examples to ensure a consistent understanding of these criteria (Appendix B).
@@ 2.2 Source Datasets
Acorn contains ratings for a diverse set of existing, newly-collected, and generated explanations. Our choice covers two commonsense reasoning benchmarks, and their respective explanation datasets of varying qualities (Figure 2). From each, we selected a random subset of 500 explanations for rating, as well as an additional 500 samples of fluency-improved versions, resulting in a total of 3,500 explanations. With five raters and eight criteria per sample, this amounts to 140k ratings in total.
Specifically, as the target commonsense reasoning benchmarks, we selected BCOPA (Kavumba et al., 2019) [^2] Balanced COPA; a superset of COPA (Gordon et al., 2012) with added “mirrored” questions that flip the correct label, i.e., the originally incorrect choice becomes correct.and CommonsenseQA (Talmor et al., 2019) based on the simplicity of their tasks and availability of large-scale explanation datasets (Wiegreffe and Marasović, 2021). Below are the respective datasets we used to source candidate explanations.
CoS-E
(Rajani et al., 2019) A widely-used explanation dataset for CommonsenseQA, albeit notoriously uninformative to humans (Nauta et al., 2023). A subset is processed through GPT-3.5333A now deprecated version of text-davinci-003. for fluency improvement, providing negative samples for higher-level issues like lack of contrastiveness or adequate information. (500 random samples + 250 fluency-improved versions)
ECQA
(Aggarwal et al., 2021) An improved version of explanations for CommonsenseQA, aligning with our criteria for well-formed explanations. (500 random samples)
Generated explanations for CommonsenseQA.
Additional high-quality explanations generated by prompting GPT-3.5 to solve a subset of CommonsenseQA, though potential issues like irrelevant information were noted. (500 samples)
COPA-SSE
(Brassard et al., 2022) Explanations for BCOPA with a subset processed through GPT-3.5 for fluency improvement. Since COPA-SSE already contains overall quality ratings, we selected a random sample of 250 questions and used each question’s top-rated and bottom-rated explanation. (500 samples + 250 fluency-improved versions) through gpt-3.5
Crowdsourced explanations for BCOPA.
ECQA’s counterpart for BCOPA; a new set of hand-written explanations, carefully crafted for contrastiveness and thoroughness. (500 samples)
Generated explanations for BCOPA.
Similarly to CoS-E, we prompted GPT-3.5 to solve BCOPA questions. (500 samples)
@@ 2.3 Rating Collection
We crowdsourced ratings for the explanations in [Uncaptioned image]Acorn using Amazon Mechanical Turk (AMT). Each rater was required to pass a qualification test, after which they were asked to participate in trial rounds, during which we addressed several clarity issues in the guidelines. The final pool was hand-picked based on their responses, resulting in 28 participants. More details on the crowdsourcing protocol are available in Appendix B. In experiments which use aggregated labels, they were produced with a majority-vote mechanism, with ties broken using the better label.
@@ 3.1 Experimental Settings
We compared four contemporary API-enabled LLMs, namely GPT-3.5, GPT-4, PALM 2, and Gemini. Each have reported high performance in diverse tasks including text-based reasoning. We used the gpt-3.5-turbo-0613 and gpt-4-0613 snapshots of OpenAI’s GPT-3.5 and GPT-4 models (Brown et al., 2020; OpenAI, 2023), and the text-bison-001 and gemini-1.0-pro versions of Google’s PALM 2 and Gemini models (Google, 2023; Gemini Team, 2023). In all experiments, temperature is set to 0.0 with 256 output tokens at most, with all other parameters left at their default values.
@@
LLMs are notoriously oversensitive to prompt format (Wadhwa et al., 2023). For the purpose of our analyses, we explored several prompting strategies and retained the most successful one for each model, as measured by correlation with majority-voted human ratings. In short, all models worked best with verbatim prompts, which corresponds to a word-by-word copy of the guidelines given to humans. The models responded with a structured list of criteria and their assigned labels for the given target explanation (in most cases). See Appendix C for a detailed overview of alternative strategies that were considered.
@@
Using free-text generation models for a classification task introduces the problem of extracting said ratings, and presents an information extraction challenge in itself. This phenomenon, inherent to generative approaches (Wadhwa et al., 2023), is a source of additional noise that affects all evaluation pipelines necessitating a non-trivial solution in real applications. In our experiments, we used a rule-based extraction method, backed up with LLM-based extraction in case of failure. Finally, we manually confirmed the remaining failures, [^4] Due to complete failure to follow the given task, exceeding the output token limit with overly verbose responses, refusal to answer due to “a lack of explanation,” or triggering content filters. and counted them as incorrect predictions. With verbatim prompts, the extraction failure rate was <0.1% for all models but Gemini (1.74%).
@@ 3.2 Inter-annotator Agreement
In subjective tasks, some degree of label variance is expected, resulting in lower inter-annotator agreement. This disagreement is not necessarily noise, but a feature of the data, as it can reflect the diversity of human opinions (Aroyo and Welty, 2015). In this context, regardless of absolute agreement, a successful LLM-based rater should be harmonious with the range of human labels rather than deviate from it.
@@ Results.
In Figure 3, the shaded area shows the agreement between human raters, while the bars show the agreement when the respective LLM’s ratings replace a random human rater. All values are averaged over twenty iterations. In most cases, agreement decreased, indicating that the LLMs generated labels that conflicted with human ones, and therefore cannot be considered suitable replacements. A notable exception is GPT-4, which maintained or improved agreement in all criteria but well-written and unnecessary information.
@@ 3.3 LLMs As A Replacement for Human Evaluation
GPT-4 maintained or improved inter-annotator in most criteria, suggesting that it does not deviate from an expected range of human ratings. Here, we ask if it can then be used as a replacement for human evaluation. Specifically, we investigated the degree to which GPT-4’s predictions align with majority-voted human labels, measured with Spearman’s rank correlation.
Figure 4: Spearman’s ranking correlation between aggregated human labels and LLM ratings.
@@ Results.
Figure 4 shows Spearman’s ranking correlation (*100 for legibility) between aggregated human labels and LLM predictions. GPT-4 results are annotated, with others shown for comparison. All values are positive, indicating a positive correlation with the ratings. The correlations in supports and contrastive were particularly strong, with values of 0.90 and 0.95, respectively. Others ranged from 0.52 to 0.78, indicating a moderately high correlation. However, the model is clearly not a perfect replacement for human raters. Instead, considering the small change in inter-annotator agreement (§3.2), the model could potentially be used as an additional rater when human annotation is scarce, which we explore in the next section.
@@ 3.4 LLMs As An Additional Rater
The results so far hinted towards GPT-4 acting similarly to an average human rater. Thus, it may seem appealing as an additional data point when human raters are scarce or expensive. Here, we verified this potential by measuring whether using GPT-4 as an additional rater improved the outcome over having fewer human raters, i.e., whether the majority-voted labels became more aligned with the original ones when the LLM is added as a rater.
Specifically, we compared Spearman’s rank correlation between the majority-voted labels with all available raters and in two alternative scenarios: one where GPT-4 is added as an additional rater and one where it is not. If the correlation increases when GPT-4 is added, it suggests that its predictions are in line with the original majority-voted labels, and it is useful as an additional rater. Otherwise, it would indicate a harmful or negligible effect, and thus its inclusion should be avoided. We repeated this comparison for scenarios with four, three, and two random human raters, respectively. All other models are also included for comparison.

### v2

@@ Abstract
Evaluating the quality of free-text explanations is a multifaceted, subjective, and labor-intensive task. Large language models (LLMs) present an appealing alternative due to their potential for consistency, scalability, and cost-efficiency. In this work, we present [Uncaptioned image]Acorn, a new dataset of 3,500 free-text explanations and aspect-wise quality ratings, and use it to evaluate how LLMs rate explanations. We observed that larger models outputted labels that maintained or increased the inter-annotator agreement, suggesting that they are within the expected variance between human raters. However, their correlation with majority-voted human ratings varied across different quality aspects, indicating that they are not a complete replacement. In turn, using LLMs as a supplement to a smaller group of human raters in some cases improved the correlation with the original majority labels. However, the effect was limited to cases where human raters were scarce, and an additional human rater had a more pronounced effect in all cases. Overall, we recommend against using LLMs as a complete replacement for human raters but encourage using them in configurations that end with targeted human involvement.
a-brassard/ACORN
@@ 1 Introduction
Natural language processing systems that not only generate correct output, but also provide an explanation (Miller, 2019) of why that particular output is correct, are desirable for several reasons, such as increasing trustworthiness (Floridi, 2019), compliance with “right to explanation” laws (e.g., European Parliament & Council of the European Union, ), increasing interpretability (Jacovi & Goldberg, 2020, but cf. Lipton, 2018 on the caveats of post-hoc explanations), as well as system improvement and knowledge discovery (Adadi & Berrada, 2018). However, this immediately raises the question of how to evaluate the plausibility of system-generated explanations in an efficient and effective manner.
Since explanations are typically free-form text, automatic evaluation of explanations suffers the well-known, but as of yet unresolved, weaknesses of automatic evaluation measures (Celikyilmaz et al., 2021), while human evaluation is characterized by low scalability, high costs, subjectivity, and inconsistency (Hartmann & Sonntag, 2022). LLM-based evaluation presents an appealing alternative due to its potential high scalability, low cost, and consistency.
@@
Specifically, we first considered whether LLMs’ labels deviate from what would be expected from human raters. We compared inter-annotator agreement between all-human annotation (five raters) and when one of the raters is replaced by an LLM. We found that stronger models maintained or marginally increased inter-annotator agreement in most cases, suggesting that it may have potential as a replacement for human judgments.
To verify this, we considered two scenarios: one where the LLM replaces the full human evaluation, and one where it is used only as an additional rater. In other words, we compared LLMs to humans collectively and individually.
With the best-performing model, Spearman’s rank correlation between majority-voted gold labels and LLM-generated ones ranged between 0.54 and 0.93, depending on the aspect, averaging 0.72. This indicates that the LLM’s labels are not entirely in line with human judgments, but they are not entirely random either, and their usefulness may depend on the specific evaluation criteria and usage case.
As for using the LLM as a single rater, we considered whether, in a limited-budget scenario, it would be beneficial to use an LLM as an additional rater. Specifically, we compared whether the majority-voted labels of a reduced set of raters and with an added LLM had a higher correlation with the original majority-voted labels. In cases with three or more raters, the addition of LLMs either had no effect or was detrimental. In cases with one or two raters, the addition of LLMs marginally improved the correlation with the original majority-voted labels, however, still less so than the addition of a human rater.
In summary, we quantified the consequences of using LLMs as a replacement or addition to human evaluation. We conclude that they are an imperfect approximation of human majority votes but still within the expected variance that comes with a subjective task, with some potential benefit when human raters are scarce. Thus, we recommend against using LLMs as a complete replacement for human raters, but encourage using them in configurations that end with targeted human involvement, such as extensively using LLMs in development or filtering stages and human experts for final testing or evaluation.
@@ 2 Building [Uncaptioned image]Acorn: Evaluation Criteria and Data Sources
In a typical commonsense reasoning test, a model selects the most plausible answer for a multiple-choice question. In a predict-and-explain setting, the model additionally generates an explanation to justify the selected answer, where we encounter the challenge of evaluating these explanations. Thus, we first define a set of rating criteria (§2.1) and collect human ratings for a selection of existing and newly collected textual explanations (§2.2, §2.3). See Appendix A for more details on the dataset, including label distributions, data source-wise average ratings, and examples.
@@ 2.1 Rating Criteria
Table 1: Explanation rating criteria used in this study. The first two criteria target consistency and general explanation quality, while the rest covers specific quality aspects.
We defined a set of criteria to target surface-level, information/content-level, and structural aspects of explanations. We also included criteria to capture (un)faithfulness and an overall rating, the latter intended to implicitly capture any other aspects considered by the raters. We defined these criteria based on common practices in natural language generation evaluation (Howcroft et al., 2020), known challenges of free-text explanations (Lipton, 2018; Rawte et al., 2023), and insights from social sciences (Miller, 2019). Table 1 summarizes the criteria. The criteria largely aligns with the fine-grained analysis conducted by Wiegreffe et al. (2022).
Supports assesses which answer the explanation supports, intended to be cross-referenced with the predicted label. A mismatch between the predicted label and the supported answer indicates a lack of faithfulness, i.e., the explanation does not reflect the model’s reasoning for the label.
Overall is a holistic assessment of the explanation, capturing any potentially informative or useful aspects that we have not explicitly covered. We encouraged workers to consider this criterion independently of the other criteria, and provided general guidelines for each star rating to ensure a consistent understanding.
Well-Written is a catch-all criterion to assess the surface-level quality of the explanation, combining criteria such as fluency, coherence, and grammaticality.
Related assesses the relevance of the explanation to the question and answers.
Factual evaluates the truthfulness of the statements in the explanation, if any, regardless of their relevance.
New Information assesses the extent to which the explanation provides new information beyond the question and answers. Workers were given the choice of none for a complete lack of new information, some for a partial addition, sufficient for a satisfactory amount of new information, and ample for highly informative explanations.
Unnecessary Information assesses the extent to which the explanation includes irrelevant information. We included this criterion to capture the challenge of generating minimal explanations.
Contrastive assesses whether the explanation contrasts the correct answer with the predicted answer.
@@ 2.2 Source Datasets
Acorn contains ratings for a diverse set of existing, newly-collected, and generated explanations. Our choice covers two commonsense reasoning benchmarks and their respective explanation datasets (Figure 2). From each, we selected a random subset of 500 explanations for rating, as well as an additional 500 samples of fluency-improved versions, resulting in a total of 3,500 explanations. The fluency-improved subset is included to prevent fluency from becoming a superficial signal, since well-written explanations typically also have high scores in all other aspects. With five raters and eight criteria per sample, this amounts to 140k ratings in total.
Specifically, as the target commonsense reasoning benchmarks, we selected BCOPA (Kavumba et al., 2019) [^1] Balanced COPA; a superset of COPA (Gordon et al., 2012) with added “mirrored” questions that flip the correct label, i.e., the originally incorrect choice becomes correct. and CommonsenseQA (Talmor et al., 2019) based on the simplicity of their tasks and availability of large-scale explanation datasets (Wiegreffe & Marasović, 2021). Below are the respective datasets we used to source candidate explanations.
CoS-E (Rajani et al., 2019) A widely-used explanation dataset for CommonsenseQA, albeit notoriously uninformative to humans (Nauta et al., 2023). A subset is processed through GPT-3.5222text-davinci-003; The model was instructed to only improve the fluency and was not given any additional context that may encourage improving the content, e.g., by supplementing related information. for fluency improvement (500 samples + 250 fluency-improved versions)
ECQA (Aggarwal et al., 2021) An improved version of explanations for CommonsenseQA, aligning with our criteria for well-formed explanations. (500 samples)
Generated explanations for CommonsenseQA. Additional high-quality explanations generated by prompting GPT-3.5 to solve a subset of CommonsenseQA, though potential issues like irrelevant information were noted. (500 samples)
COPA-SSE (Brassard et al., 2022) Explanations for BCOPA with a subset processed through GPT-3.5 for fluency improvement. Since COPA-SSE already contains overall quality ratings, we selected a random sample of 250 questions and used each question’s top-rated and bottom-rated explanation. (500 samples + 250 fluency-improved versions)
Crowdsourced explanations for BCOPA. ECQA’s counterpart for BCOPA; a new set of hand-written explanations, carefully crafted for contrastiveness and thoroughness. (500 samples)
Generated explanations for BCOPA. Similarly to CoS-E, we prompted GPT-3.5 to solve BCOPA questions. (500 samples)
@@ 2.3 Rating Collection
We crowdsourced ratings for the explanations in [Uncaptioned image]Acorn using Amazon Mechanical Turk (AMT). Each rater was required to pass a qualification test, after which they were asked to participate in trial rounds, during which we addressed several clarity issues in the guidelines. The final pool was hand-picked based on their responses, resulting in 28 participants. See Appendix B for more details on the rating collection process. In experiments which use aggregated labels, they were produced with a majority-vote mechanism, with ties broken using the better label.
@@ 3.1 Experimental Settings
We compared four contemporary API-enabled LLMs, namely GPT-4o (OpenAI, 2024), Llama-3.1 (405B) (Team, 2024), Gemma-2 (27B) (Gemma Team, 2024), and Mixtral (8x22B) (Jiang et al., 2024). Each have reported high performance in diverse tasks including text-based reasoning and represent the sate-of-the-art in generalist LLMs at the time of writing. Specifically, we used the following model versions: gpt-4o-2024-05-13, Meta-Llama-3.1-405B-Instruct-Turbo, gemma-2-27b-it, and Mixtral-8x22B-Instruct-v0.1.333We also provide the results of several smaller or older alternatives in the appendix for comparison (Tables 7 and 8). Temperature is set to 0.0 with all other parameters left at their default values.
@@
LLMs are notoriously oversensitive to prompt format (Wadhwa et al., 2023). For the purpose of our analyses, we explored several prompting strategies (listed in Appendix C) and selected the most successful one as measured by correlation with majority-voted human ratings. Most models worked best with verbatim prompts corresponding to a word-by-word copy of the guidelines given to humans, to which they responded with a structured list of criteria and their assigned labels for the given target explanation.
@@
Using free-text generation models for a classification task introduces the problem of extracting said ratings, and presents an information extraction challenge in itself. This phenomenon, inherent to generative approaches (Wadhwa et al., 2023), is a source of additional noise that affects all evaluation pipelines necessitating a non-trivial solution in real applications. In our experiments, we used a rule-based extraction method backed up with LLM-based extraction in case of failure. Finally, we manually inspected the remaining failures, [^4] Mostly due to non-compliance to the task format, e.g., responding verbosely with new labels instead of following the given choice: ”… Related: Somewhat. …” (instead of Yes or No) and excluded them from our experiments to maintain a fair comparison. The final extraction failure rates were <0.2% for all models but Gemma-2 (2.7%).
@@ 3.2 Inter-annotator Agreement
In subjective tasks some degree of label variance is expected, resulting in lower inter-annotator agreement. This disagreement is not necessarily noise but a feature of the data, as it can reflect the diversity of human opinions (Aroyo & Welty, 2015). In this context, regardless of absolute agreement, a successful LLM-based rater should be harmonious with the range of human labels rather than deviate from it.
@@ Results.
In Figure 3, the shaded area shows the agreement between human raters, while the bars show the agreement when the respective LLM’s ratings replace a random human rater ($*100$ for legibility). All values are averaged over twenty iterations. Mixtral, GPT-4o, and Llama-3.1 maintained or improved agreement in most cases, with slight decreases in supports with Mixtral, related with Llama-3.1, and with unnecessary information with GPT-4o and Llama-3.1. Gemma-2, on the other hand, decreased agreement in all but three criteria. However, the latter is also significantly smaller than the others, and illustrates the trade-off between model size and performance. Interested readers may refer to Table 7 in the appendix for all results, including several older or smaller models for comparison.
@@ 3.3 LLMs As A Replacement for Human Evaluation
The larger models maintained or improved inter-annotator in most criteria, suggesting that they do not deviate from an expected range of human ratings. Here, we ask whether they can then replace human evaluation. Specifically, we measured the degree to which the models’ predictions align with majority-voted human labels.
Figure 4: Spearman’s ranking correlation between majority-voted human labels and LLM-generated ratings ($*100$ for legibility).
@@ Results.
Figure 4 shows Spearman’s rank correlation ($*100$ for legibility) between aggregated human labels and LLM predictions. The highest values are annotated for each criterion. A comprehensive table with all results, including several smaller or older models, is available in the appendix (Table 8).
The correlation in supports and contrastive was particularly strong: 0.84 and 0.93, respectively. The unnecessary information criterion, however, stands out with a much lower correlation in all models (0.54 and less). Others ranged from 0.62 to 0.76, indicating a moderately high correlation. Overall, GPT-4o was the best-performing model in five out of seven criteria, with an average correlation of 0.72. Mixtral, the second-best model, followed closely, particularly outperforming GPT-4o in unnecessary information and contrastive.
From these results, we conclude that the larger models are relatively well-aligned with humans and could be considered effective depending on the use case. However, they are still clearly not a perfect replacement for human raters. Instead, considering the small change in inter-annotator agreement (§3.2), the model could potentially be used as an additional rater when human annotation is scarce, which we explore in the next section.
@@ 3.4 LLMs As An Additional Rater
The results so far hinted towards LLMs acting similarly to an average human rater. Thus, it may seem appealing as an additional data point when human raters are scarce or expensive. Here, we verified this potential by measuring whether using each model as an additional rater improved the outcome over having fewer human raters, i.e., whether the majority-voted labels became more aligned with the original ones when the LLM was added as a rater.
Specifically, we compared Spearman’s rank correlation between the majority-voted labels with all available raters and in two alternative scenarios: one where the model is added as an additional rater and one where it is not. If the correlation increases when the model is added, it suggests that its predictions are in line with the original majority-voted labels, and it is useful as an additional rater. Otherwise, it would indicate a harmful or negligible effect, and thus its inclusion should be avoided. We repeated this comparison for scenarios with four, three, two, and one randomly selected human rater per sample.